OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 Answer (-1, -3) (-3, 1) (-3, -1) (3, 1) Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 Answer (-1, -3) (-3, 1) (-3, -1) (3, 1) @Mathematics

myininaya (myininaya):

add (-2)*1st equation to 2nd equation

OpenStudy (anonymous):

Note 1: substitution method means solving one variable in one equation in terms of the other variable and then substituting it into the other equations. Step 1: determine which equation is easier to solve for, and solve the first equation has a coefficient of 1 before the x, so it works better. we then take the x-4y=1 and solve for x in terms of y (we could also do the other way around, but we chose this because of the 1 coefficient so we solve for x) x-4y=1 x=1+4y Step 2: substitute this into the second equation and solve for y 2x-9y=3 2(1+4y)-9y=3 2+8y-9y=3 2-y=3 y=-1 Step 3: substitute this value of y back into x in the first equation x-4(-1)=1 x+4=1 x=-3 so our solution is (-3,-1)

OpenStudy (anonymous):

omggggg xD you really tookk the time to help me out that much thank you so much p-4 :)

myininaya (myininaya):

oh it does say substitution elimination is funner so i will show you that way \[(-2x+8y)+(2x-9y)=-2(1)+3\] \[-y=1 => y=-1\] \[x-4(-1)=1 => x+4=1 => x=-3\]